# Co-VisiON: Co-Visibility ReasONing on Sparse Image Sets of Indoor Scenes

### v1

@@ Abstract
Humans exhibit a remarkable ability to recognize co-visibility—the overlapping regions visible in multiple images—even when these images are sparsely distributed across a complex scene. This capability is foundational in 3D vision and robotic perception. Despite significant progress in vision learning, it remains unclear whether current vision models have reached human-level proficiency in co-visibility analysis. In this work, we introduce the Co-Visibility reasONing (Co-VisiON) benchmark, designed to directly evaluate co-visibility reasoning on sparse image sets across over 1000 indoor scenarios. Our experiments reveal that while co-visibility is typically treated as a low-level feature matching task, it poses a significant challenge for existing vision models under sparse conditions. Notably, a proprietary vision-language model outperforms all purely vision-based approaches, with all models lagging substantially behind human performance. This gap underscores the need for more than basic pairwise vision processing—it calls for a comprehensive spatial understanding through high-level reasoning across multiple views. Inspired by human visual cognition, we propose a novel multi-view baseline, Covis, which achieves top performance among pure vision models and narrows the gap to the proprietary VLM. We hope our benchmark and findings will spur further advancements in developing vision models capable of robust, high-level reasoning in challenging, sparse environments. Our dataset and source code can be found at https://ai4ce.github.io/CoVISION.
@@ 1 Introduction
Humans are remarkably adept at constructing cognitive topology maps from multi-view observations, enabling them to accurately infer a scene’s layout. For example, when browsing websites for real estate listings or hotel bookings, one is often presented with only a few spatially sparse images. Despite the limited visual data, humans can intuitively grasp the spatial context and deduce the relationships between images, effectively constructing a mental graph where images serve as nodes and co-visible regions form the connecting edges.
Figure 1: Co-VisiON overview. If two images share non-zero co-visible surface areas, they should be connected in the binary co-visibility graph. Note that the bird’s-eye view is only for visualization and is not actually used in Co-VisiON.
@@
Modern vision models have achieved remarkable progress in these tasks and have been proficient in matching co-visible images when they have dense overlaps. However, we are particularly intrigued by humans’ robust co-visibility reasoning in sparsely distributed images, where shared visible information is limited or highly distorted. Can current vision models compete with human proficiency in co-visibility reasoning under sparse conditions?
@@
In this work, we propose the Co-VisiON benchmark, which evaluates models on Co-Visibility reasONing from sparsely captured images across over 1,000 indoor scenarios. We benchmark a diverse set of relevant methods, including feature matching [33, 18, 44, 53], contrastive learning [10], visual place recognition [4], 3D reconstruction [58, 55], supervised graph edge prediction [15, 21, 50] and large vision-language models [39, 64]. However, our results indicate that these methods fail to capture the broader context and spatial relationships within a scene.
As shown in Fig. 2, images with minimal overlap and sparse views present a significant challenge. This highlights the need for more holistic scene understanding beyond local feature matching. Beyond its challenges, co-visibility reasoning is a fundamental capability for many 3D vision applications. Having a reliable co-visibility graph can significantly improve the efficiency of structure-from-motion (SfM) [65] and simultaneous localization and mapping (SLAM) [37] under sparse observations. Similarly, it can benefit robotics tasks such as place recognition, navigation, and pose estimation, where reasoning about spatial relationships from limited viewpoints is crucial [24].
Besides the aforementioned baselines, we propose our baseline, Covis, a supervised binary classification and mask segmentation method with either pairwise or multiple views as input.
We notice that the multi-view approach, capturing more scene information, outperforms pairwise methods, provided that the additional views did not contain irrelevant information (i.e., non-co-visible features). Inspired by the confidence-level in DUSt3R [58], we propose a learnable mask supervised by co-visible regions. Applying this mask to the features filters out irrelevant information and improves co-visibility reasoning by focusing on the most relevant features.
@@
  • We propose co-visibility reasoning as a novel vision task to assess an intelligent agent’s spatial reasoning ability given sparse observations, and generate a sparse indoor image dataset with dense co-visibility annotations from photorealistic indoor simulation environments.
  • We benchmark a diverse range of baselines (including human) and demonstrate the challenge and potential of advancing co-visibility reasoning research.
  • We demonstrate that in supervised co-visibility reasoning, the multi-view method outperforms two-view ones when assisted with co-visible mask segmentation, indicating the importance of multi-view analysis in this task.
@@ 2.1 Related tasks
Visual Place Recognition. Visual place recognition (VPR) involves identifying previously seen locations and is widely studied in robotics and autonomous vehicles [34, 47, 12, 9].
The prominent VPR method, NetVLAD [4], treats VPR as an image retrieval problem, representing image features as a bag of visual words. The key distinction between VPR and co-visibility reasoning is in spatial requirements: VPR requires geographically close images, while Co-VisiON links images by overlapping surfaces regardless of distance or camera pose. Our experiments show that VPR-based methods like NetVLAD are insufficient for Co-VisiON.
Visual Localization and 3D Reconstruction. Visual SLAM simultaneously performs image-based localization and 3D mapping from (at least) RGB cameras [5, 16, 27, 36, 35, 51], whereas Structure-from-Motion [65, 46, 40] fully reconstructs 3D scenes upto a scale and estimates camera parameters in an offline fashion. In contrast, Co-VisiON builds a co-visibility graph by linking images of a scene based on shared surface visibility, without requiring full scene reconstruction or camera pose estimation.
Visual Correspondence. Estimating visual correspondence between images has wide applications in localization [37], image matching [44], and scene understanding [19]. While traditional methods focus on establishing geometric correspondence such as points [44, 41, 24], lines [65], and planes [2, 29], recent approaches focus on semantic correspondence, such as object relationships [19, 20, 17]. Notably, Rau et al. [43] proposes an image embedding method for estimating visual surface overlap—a task related to ours but limited to image pairs. However, this method performs poorly on the Co-VisiON benchmark (only 20% graph IoU). Overall, such visual correspondence tasks operate on image pairs and rely on geometric metrics, making it fundamentally different from Co-VisiON that reasons over sparse image sets with significant pose variation and is evaluated using graph IoU.
Related datasets. Benchmarking Co-VisiON requires sparse image sets with ground truth co-visibility graphs. However, many existing real-world datasets [13, 66, 30, 43, 49], including those used in [43], either do not provide such graphs or make it difficult to compute co-visible regions accurately, as those derived from point clouds or depth data are often too coarse and noisy. To obtain high-quality co-visibility ground truth, a natural solution is to leverage virtual 3D scenes with textured meshes, which can be rendered from arbitrary viewpoints with pixel-level co-visibility via OpenGL. Using photorealistic simulation environments like Gibson [62] and HM3D [42], the Co-VisiON dataset provides a sparse yet well-distributed set of viewpoints with full scene coverage and accurate co-visible correspondences, making it well-suited for advancing co-visibility reasoning.
@@ 2.2 Related methods
Feature matching. Feature matching approaches range from using handcrafted descriptors like SIFT [33] with RANSAC [18] to more recent deep-learning methods like [53, 44]. SuperGlue [44] employs a graph neural network to establish correspondences by solving an optimal transport problem on deep-learning based descriptors like SuperPoint [14]. In this work, we show such naive pairwise matching methods are not appropriate to solve Co-VisiON.
Image Representation Learning. Image representations can be learned through contrastive objectives, pairwise supervision, or spatially guided pretraining. Contrastive methods [38, 10, 22, 61, 56] bring similar images closer while pushing dissimilar ones apart. NetVLAD [4] extends this by aggregating local features into global descriptors. Pairwise classification methods predict whether two images depict the same place, typically using different backbones [15, 50, 21, 60]. While effective in simple settings, these approaches operate on image pairs and miss broader context. Alligat0R [32], as an example of spatially guided pretraining, and Covis both use co-visibility signals, but for different goals: Alligat0R uses them for pose regression, while we use them to mask irrelevant features in multi-view co-visibility prediction. The approaches above were not designed for scene-level reasoning, and our findings suggest they struggle to capture the spatial understanding.
Reasoning with vision and language models. Notably, both image matching and representation learning focus on correspondences between images from similar viewpoints or the same image with augmentations. In contrast, co-visibility reasoning requires matching images from significantly different viewpoints, demanding spatial understanding beyond feature comparison. Since VLMs [7, 39] demonstrate reasoning abilities [28, 8], we explore their potential by prompting GPT-4o [39] and Qwen-72B [1] for co-visibility reasoning between multiple images.
Multi-view learning. Multi-view learning [67] explores how information from multiple images can be integrated for tasks like object-level classification [48, 52] and 3D reconstruction [55]. Notably, MV-DUSt3R [55] fuses multi-view images to generate accurate 3D geometry with an emphasis on metric-level detail. In contrast, few works have explored multi-view learning for high-level scene understanding, which involves inferring the topological structure of a scene from sparse observations. Co-VisiON tackles this underexplored setting by jointly reasoning over multiple images to construct a co-visibility graph.
3 Co-visibility Benchmark
@@ 3.1 Task Definition
Given a sparse set of images =$\mathcal{I}=\{I_{1},I_{2},\ldots,I_{n}\}$ from a scene representing different views of a house, our objective is to construct a co-visibility graph $\mathcal{G}$. In the graph, images serve as nodes and they are connected by an edge only if they have co-visibility. Specifically, for each pair of images $(I_{i},I_{j})$, an algorithm outputs $d_{ij}$ that quantifies the degree of co-visibility between the two images.
Next, we employ thresholding to determine if $d_{ij}$ surpasses a predefined threshold $\tau$. If $d_{ij}>\tau$, we connect $I_{i}$ and $I_{j}$ with an edge in the binary graph $\mathcal{G}$. Covering all possible pairs of images, we construct an adjacency matrix $\mathbf{A}$, where each entry $a_{ij}$ is binarized from $d_{ij}$ by the threshold:
@@ 3.2 Evaluation metrics
Graph Intersection over Union (Graph IoU): Graph Intersection over Union (Graph IoU) measures the similarity between $\hat{\mathcal{G}}$ and $\mathcal{G}$ by comparing the overlap of their edges [68]. It is defined as the ratio of common edges to the total number of unique edges in both graphs:
Graph $\text{Graph IoU}=\frac{|\hat{\mathbf{A}}\cap\mathbf{A}|}{|\hat{\mathbf{A}}\cup\mathbf{A}|}.$ (2)
Area Under Curve (AUC): Since graph IoU scores can be sensitive to the chosen thresholds, we further report the Area Under Curve (AUC) to ensure fairness and robustness in task evaluation. AUC is obtained by plotting the graph IoU for different thresholds sampled evenly from the range $[0,1]$. The motivation for using AUC is that a good co-visibility reasoning method should not just achieve good IoU for a specific threshold, but also achieve consistently high IoU scores on a wide range of thresholds. A higher AUC indicates that the model is consistently better and more robust. This will particularly benefit real-world applications where threshold optimization may not be feasible. The evaluation metrics are detailed in the supplementary.
@@ 3.3 Dataset Generation
To construct a structured multi-view indoor dataset with well-defined co-visibility relationships, we employ a progressive candidate selection strategy for camera placement. Each image is selected to contribute to the overall scene representation. The selection is based on the highest intersection over union (IoU) with previously chosen images, ensuring the IoU is neither too high ($\geq$30%, reducing sparsity) nor too low ($\leq$5%, leading to a disconnected graph). This iterative process continues until a predefined coverage ratio is achieved. The resulting dataset consists of a sparsely distributed image set, with a co-visibility graph encoding pairwise relationships and corresponding co-visible region information (detailed in Sec. 3.1). We use Habitat-sim [54, 45] to collect sparse yet representative data from Gibson [62] and HM3D [42], with scene-level splits to prevent data leakage.  Tab. 1 presents dataset statistics, and Fig. I illustrates the full dataset pipeline, with additional details provided in the supplementary material.
Gibson. Gibson provides a collection of 3D models of real-world indoor environments, encompassing various types such as homes and offices. We select scenes based on the availability of .glb files and necessary metadata. Depending on the scene’s scale, the number of images ranges from a dozen to over one hundred. On average, approximately 14 images are captured per scenario for each of the 85 distinct scenes, resulting in a comprehensive dataset used for co-visibility graph annotation. The Gibson dataset is split into 80% for training and 20% for testing.
HM3D. We generate the second half of the Co-VisiON dataset from HM3D. A key difference between Gibson and HM3D is the availability of floor height data required for scene construction. While Gibson encodes floor heights in its metadata, HM3D lacks this information. To address this, we use navigable regions to distinguish floor heights and define the scene’s vertical structure. The HM3D dataset is divided into 90% for training and 10% for testing.
@@
Human Annotations. We collect human annotation as part of the Co-Vision dataset, serving as a baseline in Co-VisiON benchmark. A binary graph is generated by several experienced annotators for each scenario in Gibson. It is then compared to the ground truth to compute the graph IoU. This process represents the level of human reasoning capability for Co-VisiON.
@@ 4 Experiments on Co-VisiON
Figure 3: Multi-view Covis pipeline. Multi-view Covis is built on the MV-DUSt3R [55] encoder-decoder backbone, following the same attention pattern across views. Visual tokens for the reference, positive, and negative views are shown in blue, green, and orange respectively. Gray and bold arrows indicate cross-view attention in the decoder. The reference (anchor) uses a separate prediction head, while the positive and negative views share the classification (CLS) and mask heads. Decoder outputs are projected to pixel-wise features $F\in\mathbb{R}^{H\times W\times C}$, then filtered by a mask $M$ to obtain $F^{\prime}$.

### v3

@@ Abstract
Humans exhibit a remarkable ability to recognize co-visibility—the 3D regions simultaneously visible in multiple images—even when these images are sparsely distributed across a complex scene. This ability is foundational to 3D vision, robotic perception, and relies not only on low-level feature matching but also on high-level spatial reasoning and cognitive integration. Yet, it remains unclear whether current vision models can replicate this human-level proficiency. In this work, we introduce the Co-VisiON benchmark, designed to evaluate human-inspired co-visibility reasoning across more than 1,000 sparse-view indoor scenarios. Our results show that while co-visibility is often approached as a low-level feature-matching task, it remains challenging for existing vision models under sparse conditions. Notably, a proprietary vision-language model surpasses all vision-only baselines, but all models fall significantly short of human performance. This gap underscores the limitations of current architectures and motivates the need for models that integrate spatial and semantic information in a human-like manner. Inspired by human visual cognition, we propose a novel multi-view baseline, Covis, which achieves top performance among pure vision models and narrows the gap to the proprietary VLM. We hope our benchmark and findings will spur further advancements in developing vision models capable of robust, cognitively inspired reasoning in challenging, sparse environments. Our dataset and source code can be found at https://ai4ce.github.io/CoVISION.
@@ 1 Introduction
Humans are remarkably adept at inferring scene layout from sparse multi-view observations by integrating spatial cues and relational context. For example, when browsing websites for real estate listings or hotel bookings, one is often presented with only a few spatially sparse images. Despite the limited visual data, humans can intuitively grasp the spatial context and deduce the relationships between images, effectively constructing a mental graph where images serve as nodes and co-visible regions form the connecting edges.
Figure 1: Co-VisiON overview. If two images share non-zero co-visible surface areas, they should be connected in the co-visibility graph. Note that the bird’s-eye view is only for visualization and is not actually used in Co-VisiON.
@@
Modern vision models have shown achieved remarkable progress in these tasks and proficiency in matching co-visible images when they have dense overlaps. However, we are particularly intrigued by humans’ robust co-visibility reasoning with sparsely distributed images, where shared visible information is limited or highly distorted. Can current vision models compete with human proficiency in co-visibility reasoning under sparse conditions?
@@
In this work, we propose the Co-VisiON benchmark, which evaluates models on Co-Visibility reasONing from sparsely captured images across over 1,000 indoor scenarios. We benchmark a diverse set of relevant methods, including feature matching [32, 17, 43, 52], contrastive learning [10], visual place recognition [5], 3D reconstruction [57, 54], supervised graph edge prediction [14, 20, 49], and large vision-language models [38, 63]. However, our results indicate that these methods fail to capture the broader context and spatial relationships within a scene—highlighting the gap between current computational models and human-inspired spatial reasoning.
As shown in Fig. 2, images with minimal overlap and sparse views present a significant challenge. This highlights the need for holistic understanding beyond local features, like humans. Beyond its challenges, co-visibility reasoning is a fundamental capability for many 3D vision applications. Having a reliable co-visibility graph can significantly improve the efficiency of structure-from-motion (SfM) [64] and simultaneous localization and mapping (SLAM) [36] under sparse observations. Similarly, it can benefit robotics tasks such as place recognition, navigation, and pose estimation, where reasoning about spatial relationships from limited viewpoints is crucial [23].
Besides the aforementioned baselines, we propose Covis, a supervised binary classification and mask segmentation method with either pairwise or multiple views as input. In our experiments, we found that the multi-view approach captures more scene information and outperforms pairwise methods when irrelevant (non-co-visible) features are excluded. Inspired by DUSt3R’s confidence scores [57], we introduce a learnable mask supervised by co-visible regions to filter out irrelevant features and improve co-visibility reasoning by focusing on the most relevant features.
@@
  • We propose Co-VisiON as a novel vision task to assess an intelligent agent’s spatial reasoning ability given sparse observations, and generate a sparse indoor image dataset with dense co-visibility annotations from photorealistic indoor simulation environments.
  • We benchmark diverse baselines—including humans—and reveal the challenges and opportunities in advancing human-inspired co-visibility reasoning.
  • We demonstrate that in supervised co-visibility reasoning, the multi-view method outperforms two-view ones when assisted by co-visible mask segmentation, indicating the importance of multi-view analysis for this task.
@@ 2.1 Related tasks
Visual Place Recognition. Visual place recognition (VPR) involves identifying previously seen locations and is widely studied in robotics and autonomous vehicles [33, 46, 12, 9]. Methods like NetVLAD [5] treat VPR as image retrieval with bag-of-visual-words features. The key distinction between VPR and co-visibility reasoning is in spatial requirements: VPR requires geographically close images, while Co-VisiON links images by overlapping surfaces regardless of distance or camera pose. Our experiments show that VPR-based methods are insufficient for Co-VisiON.
Visual Localization and 3D Reconstruction. Visual SLAM simultaneously performs image-based localization and 3D mapping from (at least) RGB cameras [6, 15, 26, 35, 34, 50], whereas Structure-from-Motion [64, 45, 39] fully reconstructs 3D scenes upto a scale and estimates camera parameters in an offline fashion. In contrast, Co-VisiON builds a co-visibility graph by linking images of a scene based on shared surface visibility, not requiring full scene reconstruction or camera pose estimation.
Visual Correspondence. Visual correspondence is critical for localization [36], image matching [43], and scene understanding [18]. Traditional methods establish geometric correspondences—points [43, 40, 23], lines [64], and planes [2, 28]—while recent work explores semantic ones like object relationships [18, 19, 16]. Notably, Rau et al. [42] estimate surface overlap via image embeddings, a task related to ours but limited to image pairs and performing poorly on Co-VisiON ($20\%$ graph IoU). Overall, most correspondence methods operate on image pairs with geometric metrics, while Co-VisiON reasons over sparse sets with large pose variation and evaluates via graph IoU.
Related datasets. Benchmarking Co-VisiON requires sparse image sets with ground-truth co-visibility graphs. However, most real-world datasets [13, 65, 29, 42, 48], including those in [42], lack such graphs or produce noisy ones derived from coarse point clouds or depth maps.
To address this, we construct the Co-VisiON dataset using virtual 3D scenes with textured meshes, rendered from arbitrary viewpoints with pixel-level co-visibility via OpenGL. Built on photorealistic simulation environments like Gibson [61] and HM3D [41], it offers sparse, well-distributed views with full scene coverage and accurate co-visible correspondences, making it well-suited for evaluating co-visibility reasoning.
@@ 2.2 Related methods
Feature matching. Feature matching methods range from handcrafted descriptors like SIFT [32] with RANSAC [17] to deep-learning approaches such as LoFTR [52] and SuperGlue [43]. However, we show such naive pairwise matching methods are insufficient for Co-VisiON.
Image Representation Learning. Image representations are commonly learned through contrastive learning, pairwise supervision, or spatially guided pretraining. Contrastive methods [37, 10, 21, 60, 55], such as NetVLAD [5], aggregate local features into global descriptors. Pairwise classification approaches determine whether two images depict the same place, typically using different backbones [14, 49, 20, 59]. While effective in simple settings, these methods operate only on image pairs and fail to capture broader scene context. Alligat0R [31], as an example of spatially guided pretraining, and Covis both use co-visibility signals, but for different goals: pose regression vs. multi-view masking. These approaches were not designed for scene-level spatial reasoning and as our findings suggest, struggle to capture the spatial understanding.
Reasoning with vision and language models. Both image matching and representation learning focus on correspondences between images from similar viewpoints or the same image with augmentations. In contrast, co-visibility reasoning requires matching images from different viewpoints, demanding spatial understanding beyond feature comparison. Since VLMs demonstrate reasoning abilities [27, 8], we explore their potential by prompting GPT-4o [38], Gemini-2.0-Flash [56], SpatialRGPT [11], and Qwen-72B [1] for co-visibility reasoning between multiple images.
Multi-view learning. Multi-view learning [66] integrates information from multiple images for tasks like object-level classification [47, 51] and 3D reconstruction [54]. MV-DUSt3R [54], for example, fuses views to produce accurate 3D geometry with metric-level detail. In contrast, few works address high-level scene understanding, which involves inferring scene topology from sparse observations. Co-VisiON tackles this underexplored setting by reasoning over multiple images to construct a co-visibility graph.
3 Co-VisiON Benchmark
@@ 3.1 Task Definition
Given a sparse set of images $\mathcal{I}=\{I_{1},I_{2},\ldots,I_{n}\}$ from a scene representing different views of a house, our objective is to construct a co-visibility graph $\mathcal{G}$. In the graph, images are nodes, and they are connected by an edge only if they share co-visibility. Specifically, for each pair of images $(I_{i},I_{j})$, an algorithm outputs $d_{ij}$ that quantifies the degree of co-visibility between the two images.
Next, we employ thresholding to determine if $d_{ij}$ surpasses a predefined threshold $\tau$. If $d_{ij}>\tau$, we connect $I_{i}$ and $I_{j}$ with an edge in the graph $\mathcal{G}$. Covering all possible pairs of images, we construct an adjacency matrix $\mathbf{A}$, where each entry $a_{ij}$ is binarized from $d_{ij}$ by the threshold:
@@ 3.2 Evaluation metrics
Graph Intersection over Union (Graph IoU): This metric measures the similarity between $\hat{\mathcal{G}}$ and $\mathcal{G}$ by comparing the overlap of their edges [67]. It is defined as the ratio of common edges to the total number of unique edges in both graphs: $\text{Graph IoU}=\frac{|\hat{\mathbf{A}}\cap\mathbf{A}|}{|\hat{\mathbf{A}}\cup\mathbf{A}|}$.
Area Under Curve (AUC): AUC is computed by averaging Graph IoU across evenly sampled thresholds in $[0,1]$. AUC highlights fairness and captures performance robustness, where higher values indicate a more reliable co-visibility reasoning method, without relying on threshold tuning. Additional details are provided in the appendix.
@@ 3.3 Dataset Generation
To construct a structured multi-view indoor dataset with well-defined co-visibility, we employ a progressive camera placement strategy. Each image is selected to best complement scene coverage, based on its intersection over union (IoU) with previously chosen views. To ensure image diversity and graph connectivity, IoU is constrained between 5% and 30%. This iterative process continues until a target coverage is reached. The resulting dataset consists of a sparsely distributed image set, with a co-visibility graph encoding pairwise relationships and corresponding co-visible region information (detailed in Sec. 3.1). We use Habitat-sim [53, 44] to collect sparse yet representative data from Gibson [61] and HM3D [41], with scene-level splits to prevent data leakage.  Tab. 1 presents dataset statistics, and Fig. I illustrates the full dataset pipeline, with additional details provided in the supplementary material.
Gibson. Gibson provides 3D models of real-world indoor environments, including homes, offices, and more. We select scenes based on the availability of .glb files and required metadata. Depending on the scene size, each includes between a dozen and over a hundred images. On average, 14 images are captured per scenario across 85 distinct scenes, forming a comprehensive dataset for co-visibility graph annotation. We use an 80/20 split for training and testing.
HM3D. We generate the second half of the Co-VisiON dataset from HM3D. Unlike Gibson, HM3D lacks floor height metadata. To recover vertical structure, we infer floor heights using navigable regions. The dataset is split 90% for training and 10% for testing.
@@
Human Annotations. We collect human annotations in the Co-VisiON dataset as a benchmark baseline. For each Gibson scene, several experienced annotators generate a co-visibility graph, which is then compared to the ground truth to compute graph IoU. This provides an estimate of human-level performance on the Co-VisiON task.
@@ 4 Experiments on Co-VisiON
Figure 3: Multi-view Covis pipeline. Multi-view Covis is built on the MV-DUSt3R [54] encoder-decoder backbone, following the same attention pattern across views. The encoder produces initial features $F_{0}$ for each view, and the decoder refines these features layer-by-layer, producing intermediate outputs $F_{1},\ldots,F_{d-1}$. Decoder outputs $F_{d}$ are projected to pixel-wise features $X$ via MLP, then subsequently filtered by a mask $M$ to obtain $X^{\prime}$. Visual tokens for the reference, positive, and negative views are shown in blue, green, and orange respectively. Gray and bold arrows indicate cross-view attention in the decoder. The reference (anchor) uses a separate prediction head, while the positive and negative views share the classification (CLS) and mask heads.